OpenStudy (kaylala):

Trigonometric Equation 2tan^2A + 3secA = 0

OpenStudy (kaylala):

@hartnn Another one please

hartnn (hartnn):

which pythagorean identities do you know ?

OpenStudy (kaylala):

I've bee through all except for the half identities

OpenStudy (kaylala):

@hartnn

hartnn (hartnn):

so which identity relates tan^2 A and sec^2 A ?

OpenStudy (kaylala):

Sin?

OpenStudy (kaylala):

I mean cos?

hartnn (hartnn):

ever sen this ? \(\\tan^2 A =sec^2A-1\) ? use it here

OpenStudy (kaylala):

Yes. I'm familiar with that Wait. Let me try

OpenStudy (kaylala):

I got 2x^2+3x-2 Is it right? So i may move on

hartnn (hartnn):

yes, that = 0 remember its an equation, not an expression

OpenStudy (kaylala):

Oh right

OpenStudy (kaylala):

How'll this be nothing will equal to 3x?

OpenStudy (kaylala):

@hartnn ^

hartnn (hartnn):

what ? i didn't get your question...

hartnn (hartnn):

2x^2+3x-2 = 0 x= ..... x=....

OpenStudy (kaylala):

Oh sorry What i meant was what should i do next? How do i find x? I cant use the FOIL method.

hartnn (hartnn):

factor method. 2 numbers with sum = +3 and product = -4

OpenStudy (kaylala):

Got it Wait

OpenStudy (kaylala):

Sec = 1/2 Sec={no solution Sec=-2 No solution ????? I dont know Hpw's this? @hartnn

hartnn (hartnn):

sec A can be -2

OpenStudy (kaylala):

I dont see any negative 2 in the graph list Sec A =-2

hartnn (hartnn):

sec A = -2 secA = 1/cos A cos A = -1/2 when ?

OpenStudy (kaylala):

I dont see any cos A =-1/2 There is only cos A=1/2

OpenStudy (kaylala):

Or did i get it wrong?

OpenStudy (kaylala):

@hartnn ^

hartnn (hartnn):

how about angles 120 and 210 ?

OpenStudy (kaylala):

Where did you get that. Those arent in the graph that was given to us

OpenStudy (kaylala):

Oh i see now So will the answer for secA=-2 or cosA=-1/2 be: A={120,210}

OpenStudy (kaylala):

Is that it already? How about sec=1/2? Does it really have no solution? @hartnn

hartnn (hartnn):

yes. in the range 0 to 360. range of sec function does not include values between -1 and 1 so, sec A = 1/2 will have no solution, (in the same way cos A = 2 have no solution)

OpenStudy (kaylala):

Okay thanks again

hartnn (hartnn):

welcome ^_^